OpenStudy (anonymous):

How many g of water are required to be mixed with 11.75 g of HgCl in order to make a 0.01 m solution? (Refer to the periodic table for atomic weights.) 40 g 5,000 g 2 x 106 g 2 x 108 g

OpenStudy (anonymous):

Molar mass HgCl = 200.59+35.453 = 236.043g/mo; 11.75g = 11.75/236.043 = 0.04977 = 0.05mol HgCl If you mix this with 1000g H2O you will have 0.05m solutiuon You want a 0.01m solution: mix with 0.05/0.01*1000 = 5,000g H2O